ANOMALOUS HYPERFINE STRUCTURE OF NSF_{3} IN THE DEGENERATE VIBRATIONAL STATE v_{5}=1: LIFTING OF THE PARITY DEGENERACY BY THE FLUORINE SPIN-ROTATION INTERACTION

H. HARDER, S. MACHOLL, H. MAEDER, Institut für Physikalische Chemie, Universität Kiel, 24098 Kiel, Germany; L. FUSINA, Dipartimento di Chimica Fisica e Inorganica, Università di Bologna, 40136 Bologna, Italy; \underline{I. OZIER}, Department of Physics and Astronomy, University of British Columbia, BC V6T 1Z1, Canada.

For the principal isotopomer ^{14}N^{32}S^{19}F_{3} of thiazyl trifluoride in the degenerate fundamental state (v_{5}=1), the hyperfine structure has been investigated in the Q-branch spectrum between 8 and 26.5 GHz using microwave Fourier transform waveguide spectrometers with a resolution limit of ≈ 30 kHz. In addition to l-type doubling spectra and l-type resonance transitions with (Δk =Δl=±2), perturbation-allowed spectra were measured with Δ(k − l) =±3, ±6. The range in J was from 13 to 61; for the lower states, kl=−3, −2, −1, 0, +1. For all the transitions, the hyperfine patterns observed are predicted to be doublets when only the nitrogen quadrupole Hamiltonian H_{Q}^{N} is taken into account. Doublets were indeed measured for transitions with Γ_{RV} =A_{1} ↔ A_{2}, where Γ_{RV} is the rovibrational symmetry. However, when Γ_{RV} =E ↔ E, triplets and quartets were observed in addition to doublets. These anomalous hyperfine patterns are shown to be due to the (Δk=±1) and (Δk=±2) matrix elements of the fluorine spin-rotation Hamiltonian H_{N}^{F} characterized by the fluorine spin-rotation constants c(1) = (c_{xz} +c_{zx}^{∗}) and c(2) = (c_{xx}−c_{yy}), respectively. These terms in H_{N}^{F} lift the parity degeneracy for Γ_{RV} =E. The rovibrational Hamiltonian H_{RV} was adopted from an earlier partner study.^{1} A good fit to the hyperfine data was obtained with a standard deviation of 3.1 kHz. In the fitting process, 12 rovibrational parameters were varied, while the remaining constants in H_{RV} were left at the values of Ref. (1). In addition, 6 hyperfine parameters were determined: four in H_{Q}^{N}, and two in H_{N}^{F}. It was found that |c(1)| = 7.48(24) kHz and |c(2)| = 2.423(22) kHz. This determination of |c(1)| is the first to be reported based on frequency measurements. The key to the observation of the parity doubling lies in the severe mixing into the eigenvectors of basis vectors with several different values of kl as a result of the clustering^{1} of rovibrational levels at low K.

1S. Macholl, H. Harder, H. Mäder, L. Margulès, P. Dréan, J. Cosléou, J. Demaison, and P. Pracna, J. Phys. Chem. A 113, 668